### v1

# AutoResearcher: Automating Knowledge-Grounded and Transparent Research Ideation with Multi-Agent Collaboration

Abstract.
Effective research relies on organizing extensive information and stimulating novel solutions. Agentic systems have recently emerged as a promising tool to automate literature-based ideation. However, current systems often remain black-box. Their outputs may appear plausible but weakly grounded, with limited transparency or control for researchers. Our work introduces AutoResearcher, a multi-agent demo system for knowledge-grounded and transparent ideation. Specifically, AutoResearcher integrates meticulously designed four stages into a unified framework: (A) Structured Knowledge Curation, (B) Diversified Idea Generation, (C) Multi-stage Idea Selection, and (D) Expert Panel Review & Synthesis. Different from prior pipelines, our system not only exposes intermediate reasoning states, execution logs, and tunable agents for inspections, but also enables the generation of hypotheses that are both diverse and evidence-aligned. Our design is also domain-agnostic: as long as literature sources exist, the same pipeline can be instantiated in any scientific field. As an illustrative case, we demonstrate AutoResearcher on a graph-mining case study (-truss breaking problem), where it generates distinct, plausible hypotheses with evidence and critiques. A live demo and source code are available at https://github.com/valleysprings/AutoResearcher.
1. Introduction
The formulation of novel research ideas is a central driver of scientific progress. It remains one of the most challenging and time-intensive stages of research-oriented inquiry, since effective research depends on organizing large volumes of information and cultivating original, diverse, and impactful solutions. As many fields expand at an unprecedented pace, researchers confront severe information overload: the literature has grown beyond what any individual can reasonably process. At the same time, cognitive constraints, such as bias, fixation, and narrow search strategies, further limit the exploration of genuinely novel hypotheses. This tension between accelerating knowledge production and bounded human attention creates a bottleneck for innovation.
Foundation models, particularly large language models (LLMs), can rapidly reorganize knowledge at scales beyond human capacity. Emerging evidence suggests that they can foster creative and divergent thinking across scientific domains (Radensky et al., 2025; Guo et al., 2024). Yet, the challenge lies not merely in demonstrating that LLMs can generate ideas, but in structuring those ideas so they remain evidence-based, diverse, and actionable for researchers to evaluate.
Prior work motivating our design falls into two complementary strands: (1) LLM-assisted ideation, where prompting strategies and structured human input broaden candidate spaces and reduce fixation (Pu et al., 2024; Liu et al., 2024); and (2) agentic pipelines, in which multiple LLMs coordinate across complementary steps such as retrieval, planning, and knowledge-graph integration for research ideation, thereby enhancing traceability, aligning outputs with evidence, and reducing hallucinations (Li et al., 2024; Xiong et al., 2024; Yamada et al., 2025; Guo et al., 2024).
Despite rapid progress, existing automated research ideation systems still face two major challenges. First, the trade-off between novelty and grounding: expansive prompting can generate a wide range of ideas, but often without strong grounding, while constrained prompting yields more plausible outputs but suppresses exploration (Guo et al., 2024; Radensky et al., 2025). Second, robustness and transparency: current LLM workflows still face challenges in robustness and reliability, often failing unpredictably across domains or tasks. At the same time, their internal reasoning is largely opaque, making it difficult for researchers to understand, verify, or debug the generated ideas. Efforts to improve robustness and transparency exist, but typically come with significant complexity (Baek et al., 2025). Taken together, these gaps leave current research ideation systems unbalanced. While progress has been made on each front individually, a trustworthy, knowledge-grounded, and transparent system for real-world scientific applications is still lacking (Wei et al., 2025).
To address these challenges, we introduce AutoResearcher, a multi-agent framework for knowledge-grounded and transparent research ideation. Our work advances research ideation along two dimensions: system-level design and component-level designs. At the system level, we propose a compact four-stage design that mirrors the real-world use of ideation tools, improving efficiency and maintaining methodological control. Specifically, it consists of (A) Structured Knowledge Curation: Anchors the process through topic decomposition, retrieval, and KG construction, organizing evidence into a structured and traceable context. (B) Diversified Idea Generation: Transforms the grounded context into diverse yet structured candidate hypotheses via planning, decomposition, and diverse generation strategies; (C) Multi-stage Idea Selection: Combines internal scoring with external similarity checks against the retrieved literature to filter redundant or weakly-supported candidates. (D) Expert Panel Review & Synthesis: Integrates parallel peer-style reviews to consolidate selected ideas into a coherent proposal. By linking planning, idea generation, idea selection, and critique to curated structured evidence, AutoResearcher promotes hypotheses that are traceable to prior works and relevant background knowledge.
At the component level, AutoResearcher incorporates mechanisms to handle LLM limitations, including hallucination, prompt sensitivity, and uneven domain coverage. At specific steps, it integrates knowledge grounding, iterative self-refinement, and optional human-in-the-loop supervision. The grounding module aligns the model’s reasoning with verified knowledge sources to ensure factual consistency, the refinement module enables multi-round critique and revision, and the supervision interface allows human experts to guide or validate the reasoning process when needed. Different from prior idea generators that function as black-box LLM pipelines, AutoResearcher is designed as an interactive and traceable system. Users can track intermediate hypotheses, critiques of candidate ideas, and engage with the system in real time.
In summary, this work makes three key contributions:
We present AutoResearcher, a multi-agent system for knowledge-grounded and transparent research ideation.
It consists of four key components: (A) Structured Knowledge Curation, (B) Diversified Idea Generation, (C) Multi-stage Idea Selection, and (D) Expert Panel Review & Synthesis. Throughout these stages, it integrates knowledge grounding, iterative self-refinement, and optional human-in-the-loop supervision for enhanced performance.
We release a web demonstration showcasing the interactive workflow and outputs of the system on the -truss discovery use case, illustrating how LLM-based agents can support real-world scientific research ideation.
2. System Design
Figure 1 presents the architecture of AutoResearcher, which integrates four core modules: structured knowledge curation, diversified idea generation, idea selection, and expert panel review. Each module mirrors a corresponding phase in human research iteration, from active grounding from prior research to generating, refining, and evaluating hypotheses. Details of each component are provided in Sections 2.1–2.4.
2.1. Structured Knowledge Curation
Like human researchers who first survey existing literature before generating new ideas, AutoResearcher begins by constructing a structured KG that organizes retrieved papers into a traceable state. This KG provides well-grounded contexts for downstream idea generation. It consists of steps of multi-granularity retrieval and incremental KG construction.
Multi-granularity Retrieval. The process starts with LLM-guided topic decomposition, expanding a seed query into multiple concepts spanning domain terminology, methodologies, and applications. Instead of relying on a single handcrafted query, AutoResearcher performs a combinatorial search at varying granularities, retrieving a diverse yet focused literature set via the Semantic Scholar API.
The process starts with an LLM-guided topic decomposition, where the seed topic is analyzed to extract key domain terms, methodological cues, and application areas. Rather than issuing a single query, AutoResearcher expands these facets into multiple query combinations of varying granularity, for example: individual keywords, pairwise combinations, and higher-order tuples, to maximize recall. Each query is executed via the Semantic Scholar API, and results are merged after deduplication using semantic filtering. This combinatorial retrieval process broadens coverage across related subtopics while maintaining topical relevance.
Incremental KG Construction. To achieve balanced coverage and interpretability, AutoResearcher constructs the KG in four controlled phases. First, LLM-based entity extraction identifies core problems, methods, and applications to form the conceptual backbone. Second, mini-batch enrichment adds entities and relations from paper metadata while maintaining contextual coherence. Third, degree-based expansion samples top- high-degree nodes (default ) to mine adjacent or emerging concepts. Finally, hybrid sampling (60% high-degree, 40% random) uncovers latent methodological and theoretical links without introducing new entities. This incremental process yields a structured and extensible KG that combines precision with exploratory breadth.
2.2. Diversified Idea Generation
Building on the curated KG, AutoResearcher generates research ideas through a hierarchical, multi-strategy process that mirrors human brainstorming, consisting of steps of literature-informed planning, graph-of-thought exploration, multi-strategy variant generation and iterative refinement. The goal is to produce a diverse yet coherent idea pool.
Literature-Informed Planning We introduce a planner agent that serves as the bridge between knowledge grounding and idea synthesis. It analyzes top- high-degree entities and summarizes the top- semantically relevant papers to extract key research cues. Through LLM-based gap analysis, it identifies open limitations and decomposes them into three aspects: Problem Statement, Proposed Methodology and Experimental Validation. This structured blueprint provides the foundation for large-scale idea generation.
Graph-of-Thought Exploration. To leverage structural grounding, we adopt Graph-of-Thought (GoT) reasoning (Besta et al., 2024) to sample KG-grounded reasoning traces. The GoT module connects facet nodes to high-degree KG entities and performs asynchronous depth-first sampling (branching factor , depth ), scoring paths by node quality (), edge-type diversity (), and length preference (). High-quality paths are retained, each encoding a distinct research trajectory that enriches idea diversity.
Multi-Strategy Variant Generation. To ensure breadth, AutoResearcher applies an over-generation factor () and executes strategies in parallel: (1) Base variants extend the faceted plan; (2) GoT variants reformulate high-scoring reasoning paths into structured proposals; and (3) Cross-pollination, triggered on demand, synthesizes hybrids of top-ranked ideas guided by KG cross-connections. Redundant ideas are pruned via real-time string and semantic matching.
Iterative Refinement. The refined pool undergoes multi-round critique and validation elaboration, expanding experimental facets (datasets, metrics, ablations, stakeholder evaluation) while preserving revision history, scores, and reasoning traces. The resulting transparent idea pool captures reasoning traces, revision history, and top- supporting literature.
2.3. Idea Selection
To ensure both coherence and novelty, AutoResearcher performs a two-stage selection.
Internal Selection. Each idea is evaluated across four weighted criteria: novelty (), feasibility (), clarity (), and impact () to generate preliminary rankings. Top ideas undergo iterative merging: pairs with Jaccard similarity are LLM-merged into unified proposals until convergence criteria are met. This process eliminates redundancy via evidence-based consolidation.
External Selection. Remaining ideas are compared to retrieved literature using BGE-M3 embeddings (Chen et al., 2024) and selected using cosine similarity over combined text fields. Candidates with maximum similarity below are retained, while the system logs top overlapping papers for transparency.
2.4. Expert Panel Review & Synthesis
Inspired by scientific peer review, AutoResearcher employs a multi-agent evaluator to ensure that generated ideas are not only diverse and novel but also feasible, rigorous, and well-justified. Two specialized agents operate asynchronously to review all selected ideas in parallel. The reviewer agent focuses on technical soundness and feasibility, while the novelty agent assesses originality and contribution relative to prior work. Both agents follow a structured scoring template derived from standard conference and journal review rubrics, evaluating each idea across five dimensions (feasibility, expected impact, technical soundness, implementation complexity, and distinctiveness from prior work) on a 1–5 scale, with qualitative feedback on strengths, weaknesses, and recommended revisions.
An aggregator module fuses both perspectives by averaging dimension-level scores (equal weighting) and consolidating textual feedback into a single meta-review. A unified score is then computed using a weighted aggregate that emphasizes feasibility and originality, with all sub-scores preserved for traceability. Ideas with unified scores above 3.5 (approximately corresponding to a “weak accept” in peer-review terms) are classified as high-quality and prioritized for inclusion. If the number of high-quality ideas exceeds the target portfolio size, all are retained to avoid discarding promising work; if fewer meet the threshold, remaining slots are filled by the next best candidates.
3. Demonstration
We illustrate the main user interface and a live case study of AutoResearcher; additional analyses and materials are provided in the Appendix and project repository.
3.1. Interface
Figure 2 presents the demonstration interface of AutoResearcher. The interface is organized into four functional regions: (a) Session control, where users define the research topic, number of ideas, and launch new sessions; (b) Status monitor, displaying the current phase, runtime progress, and metadata; (c) Final output, summarizing session results with structured downloads; and (d) Real-time logs, showing detailed traces such as phase transitions, runtimes, and agent activities. This layout emphasizes transparency and interactivity: Users can track the entire pipeline in real time, with each component’s output preserved as structured JSON files for inspection or reuse.
3.2. Case study
The live demonstration focuses on the -truss breaking problem: given an undirected graph , an integer , find the smallest edge set such that the deletion of makes contain no -truss. This task is combinatorially complex: removing one edge can trigger cascading effects, and the current research is still ongoing, which makes it an ideal stress test for AutoResearcher retrieval, reasoning, and researching capabilities.
(A) Structured Knowledge Curation. The system retrieves relevant papers on -truss decomposition and breaking (e.g., scalable decomposition algorithms, parallel index construction, and dynamic update techniques) and builds a lightweight knowledge graph to anchor the session in evidence.
(B) Diversified Idea Generation. From this curated context, it produces multiple candidate directions. In the -truss session, these include scalable localized algorithms that bypass global recomputation, epidemic containment strategies on temporal contact networks, and learning-based edge importance prediction. Each idea is expanded into a structured proposal with a problem statement, a methodological sketch, and a validation plan.
(C) Multi-stage Idea Selection. Candidate ideas are filtered by internal similarity checks and external alignment with the literature. Redundant variants are removed by LLM-based deduplication, while embedding similarity ensures topical relevance. This narrows the pool to distinct, high-quality candidates.
(D) Expert Panel Review & Synthesis. Reviewer-style agents then provide numerical scores and textual critiques. For instance, the scalable algorithm idea was rated highly for novelty (4.2) and clarity (4.5), with weaknesses noted on scalability for extremely large graphs. The portfolio analysis consolidates the evaluations, yielding five high-quality candidates (average overall score 4.1/5). The resulting portfolio spans multiple directions, including scalable algorithm design, epidemic containment in temporal networks, theoretical decomposition results, distributed intervention strategies, and learning-based approaches.
A detailed description of one representative idea generated by AutoResearcher, the Scalable Approximation Framework for Real-Time k-Truss Disruption in Massive Dynamic Graphs, is included in Appendix LABEL:app:idea for reference.
4. Conclusion
In this work, we present AutoResearcher, a compact four-stage pipeline for knowledge-grounded and transparent ideation: Structured Knowledge Curation, Diversified Idea Generation, Multi-stage Idea Selection, and Expert Panel Review & Synthesis. It generates a broad, non-redundant set of candidate ideas, expanding the researcher’s space for exploration. At the same time, intermediate states and evidence links are exposed, ensuring that ideas are verifiable and their provenance is clear.
Appendix A Agents Descriptions for AutoResearcher
AutoResearcher is an autonomous multi-agent system designed for generating, evaluating, and refining research ideas in a literature-grounded manner. It consists of multiple specialized agents organized into a 4-stage pipeline that spans from literature retrieval to final portfolio analysis. Each agent performs a clearly defined function, ensuring modularity, transparency, and reproducibility.
A.1. Primary Agents
Knowledge Grounding Agent
This agent performs academic literature retrieval and ranking using the Semantic Scholar API. It decomposes the input topic into conceptual units, generates combinatorial search queries, and retrieves the top- most relevant papers. The agent integrates citation-based weighting and semantic relevance scoring while performing asynchronous, rate-limited API calls. Output is a ranked set of literature references used for downstream reasoning.
KGBuilder Agent
KGBuilder orchestrates multi-stage KG construction driven by an LLM. It seeds the graph from a topic, iteratively expands concepts, and enriches relationships while tracking entity types. Literature ingestion batches papers, extracting authors, keywords, methodologies, and temporal metadata before linking documents through shared themes. Sampling heuristics steer expansion targets and relationship probing using degree metrics and randomness to cover high-value concepts.
IdeaGenerator Agent
The central ideation engine employs multi-strategy synthesis to generate literature-informed research ideas. Leveraging a KG constructed from retrieved papers, it produces initial planning ideas, generates diverse variants, and cross-pollinates them for novelty. Over-generation (typically the target count) and self-critique cycles enhance diversity and rigor. Internal sub-agents include the Research Planner, Facet Decomposition, and GoT Variant, which will be discussed later.
Internal Selection Agent
This agent performs early-stage idea filtering and merging. It applies pairwise or clustering-based similarity analysis (using Jaccard metrics) to consolidate overlapping ideas, followed by LLM-assisted synthesis.
External Selection Agent
Ensuring distinctness from prior work, this agent computes idea–literature similarity using multiple backends (TF-IDF, BGE, Stella, Jina embeddings). It labels overlap levels and provides actionable recommendations for improving conceptual distinctness. This stage filters out near-duplicate or low-novelty proposals before review.
Reviewer Agent
Emulating academic peer review, this agent evaluates ideas across five dimensions: novelty, feasibility, clarity, impact, and methodological soundness. It produces structured feedback with quantitative scores, bullet-point strengths and weaknesses, and final recommendations. Weighted scoring enables systematic portfolio-level comparison.
Novelty Agent
A specialized assessor focusing on originality, interdisciplinary potential, and paradigm-shift likelihood. It quantifies novelty across four sub-dimensions and classifies contributions as high, mid, or low potential. Trend analysis across ideas reveals thematic innovation patterns.
Portfolio Analysis
The Aggregator fuses evaluations from the Reviewer and Novelty agents into a unified portfolio assessment. It computes composite scores, ranks top ideas, and synthesizes common strengths, weaknesses, and improvement insights. The agent outputs a portfolio-quality report with statistical summaries and top-ranked ideas.
A.2. Sub-Agents in IdeaGenerator
Research Planner
This module transforms literature and KG context into structured research plans, decomposed into problem, methodology, and validation facets. It identifies gaps, integrates supporting evidence, and iteratively refines plans through feedback loops.
Facet Decomposition
Responsible for structuring ideas into three core components (Problem Statement, Proposed Methodology, and Experimental Validation). This module ensures completeness and coherence across facets. It also enriches ideas with KG-derived context for scientific depth.
GoT Variant
A reasoning subsystem that sample diverse paths from the KG. It samples high-quality reasoning paths via depth-controlled DFS and scores them based on node centrality, edge diversity, and path coherence. The resulting paths inform GoT-based idea variants with structured reasoning.
Cross-Pollination
The Cross-Pollination Agent asynchronously fuses high-quality candidate ideas. Each pairing is guided by cross-cluster connectors distilled from the topic-centered evolving KG, promoting conceptual transfer across domains.
Idea Exploration
The Idea Exploration Agent performs two-layered refinement passes to transform preliminary candidates into internal-ready proposals. In the first pass, it launches an asynchronous self-critique suite where each idea is evaluated and revised. The second pass conducts batched elaboration on methodology and validation aspects, guided by critique-derived signals.
Appendix B Hyperparameter of AutoResearcher
Overview.
This appendix summarizes the main hyperparameter of AutoResearcher based on four stages: A) Structured Knowledge Curation, B) Diversified Idea Generation, C) Multi-stage Idea Selection, and D) Expert Panel Review & Synthesis. Tables 1–4 provide a detailed breakdown of our parameter setting. For the criteria used, N/F/C/I/M denotes novelty, feasibility, clarity, impact, and methodology soundness.
Global LLM settings
Unless specified, we use GPT-5 / GPT-4.1 as default model generator, with temperature set at 0.8, max tokens per operation: 32K.
| Agent | Hyperparameter | Value / Options |
| Knowledge Grounding | Papers per run | 50 |
| Scoring weights | relevance 0.7, citation 0.3 | |
| KG Builder | Processed batch size | 3 |
| Seed sampling | size 10, sampled based on degree | |
| Relation enhancement | sample 20, top-degree 0.6, random 0.4 | |
| Enhancement iters | 2 |
| Agent | Hyperparameter | Value / Options |
| IdeaGenerator | Requested Ideas | 3 |
| Over-generation factor | 10 | |
| Variant generation strategies | base/GoT/cross-pollination, equal weighted | |
| Refinement Rounds | 2 (maximum) | |
| GoT | Max paths / entities | 100 / 20 |
| DFS depth / branching | 5 / 3 | |
| Start nodes | 10 | |
| Min path score | 0.5 | |
| Scoring weights | node quality 0.6, edge diversity 0.2, len 0.2 | |
| Bridges per facet | 3 |
| Agent | Hyperparameter | Value / Options |
| Internal Selection | Criteria weights | N/F/C/I = 0.30/0.25/0.20/0.25 |
| Batch size | 50 | |
| Merge similarity (Jaccard) | 0.85 (max merge iters 20) | |
| Selection multiplier | from 10x to 5x | |
| External Selection | Similarity thresholds | low <0.3, mid 0.5--0.7, high >0.7 |
| Embedding backend | BGE-M3, Stella v5, Jina v3 (optional) |
| Agent | Hyperparameter | Value / Options |
| Reviewer | Criteria weights | N/F/C/I/M = 0.25/0.20/0.20/0.25/0.10 |
| Decision thresholds | accept 4.0, minor rev. 3.5 | |
| Novelty | Criteria weights | tech./prob./app./theory/empirical, equal-weighted |
| Contribution bands | high 4.0, mid-high 3.5 | |
| Portfolio Analysis | Review/novelty suggestions | 3 / 2 |
| Good-score threshold | 4.0 | |
| Portfolio bands | excellent 4.0, good 3.5, mid 3.0, weak 2.0 |
Appendix C Prompt Templates for AutoResearcher
Here, we provide templates used for AutoResearcher. Researchers may adapt or extend these prompts for reproducibility and further experimentation. Note that only a subset of the prompts is presented in this section, while the complete collection can be accessed in the project repository.
C.1. Structured Knowledge Curation
This module defines the structured prompting framework for knowledge grounding and graph construction. It provides a systematic way to (1) decompose complex research topics into layered, searchable representations and (2) extract structured entities and relations suitable for lightweight KG construction.